Record the accounting transaction and then the audit procedures for the economic event described.  Also include a discussion of any high risk audit areas.

        

On 3/14/20, the FCOB purchased 22 Sony monitors for $1,290 each, on account.  Sony’sfiscal year end is March 31.  Included in the sale price is a free 3-year extended service warranty package valued at $105 per monitor. Sony has a 31% gross profit percentage.

1.a What is the Accounting Transaction at Date of Sale?

1.b What are the Audit Procedures?
Answer #1

1. a What is the Accounting Transaction at Date of Sale?

Basis accounting of service type warranties under IFRS 15, receipts from customers for service warranties are accounted as performance obligations on date of receipt and recognizes as revenue when performance obligation is satisfied.

Sale value per unit excluding warranty cost= 1290-105= $1185

Total value including warranty cost= 1290*22= $28,380

Total value excluding warranty cost= 1185*22= 26,070

Total warranty cost = 105* 22= $2,310

Accounting entry on transaction date

Bank A/c Dr.               28,380

            To revenue from sale of monitors      26,070

            To Contract liability                             2,310

1. b What are the Audit Procedures?

The following points to be taken into consideration when auditing revenue recognized under Service type warranties.

· Identify the customer contract(s)

· Identify the performance obligation(s) in the contract

· Determine the transaction price

· Allocate the transaction price to the performance obligation(s) in the contract

· Recognize revenue when (or as) the entity satisfies a performance obligation